# Earliest Galaxy Evolution in the CANUCS+Technicolor fields: Galaxy Properties at $z\sim 10-16$ seen with the Full NIRCam Medium and Broad Band Filters

### v1

@@ 4.3 Cosmic UV Luminosity Density
Given the parameterized UV LF are measured as a function of redshift directly from the data in Sec. 4.1, we can directly compute the redshift evolution of UV luminosity density ($\rho_{\rm UV}(z)$) from the UV LF measurement. We integrate the LF in the Schechter functional form (Eqn. 4.1) down to $M_{\rm UV}=-17$ mag, and obtain the marginalized $\rho_{\rm UV}(z)$ estimation from the MCMC sampling in Sec. 4.1. Figure 4 shows the median estimation (black solid) with the 16th-84th percentile (red shaded) from our results, comparing with the literature (P. G. Pérez-González et al., 2023, 2025; N. J. Adams et al., 2024; C. T. Donnan et al., 2024; B. Robertson et al., 2024; C. J. Willott et al., 2024; Y. Harikane et al., 2025). The overall $\rho_{\rm UV}$ value in this work is somewhat lower than most JWST studies, and in line with N. J. Adams et al. (2024) and C. J. Willott et al. (2024). Some literature claims that significant modifications of dust, IMF, or star formation efficiencies (SFEs) in model assumptions are required based on the overabundance of UV bright galaxies at $z>10$ (e.g., S. L. Finkelstein et al., 2024; Y. Harikane et al., 2025; P. G. Pérez-González et al., 2025). However, our result together with N. J. Adams et al. (2024) and C. J. Willott et al. (2024) should suggest that such significant modifications are not necessarily required.
@@
The log-linear rate of the UV LF evolution $\eta$ is directly connected to the $\rho_{\rm UV}(z)$ redshift evolution, and our result indicates $\rho_{\rm UV}(z)$ decline as $d\log(\rho_{\rm UV})/dz=-0.21^{+0.11}_{-0.12}$ at $1610<z<$ 16. This is less steep than measurements at $z\sim 8-11$ by N. J. Adams et al. (2024); C. J. Willott et al. (2024) and consistent with B. Robertson et al. (2024) at $z=11-16$. Importantly, this is also less steep than the prediction by the simple constant extrapolation of the SFE from lower redshift (e.g., Y. Harikane et al., 2022, the gray dash-dot line in Figure 4). These results suggest $\rho_{\rm UV}$ evolution gets shallower above $z\sim 11$, and one of the possible interpretations is the SFE may start to deviate at $z>11$ from the simple constant extrapolation from lower redshift.
@@
Figure 4 also compares observations to predictions by several simulations: THESAN-ZOOM (R. Kannan et al., 2025), Universe Machine (P. Behroozi et al., 2020), FLARES (A. P. Vijayan et al., 2021; S. M. Wilkins et al., 2023), and SC SAM (L. Y. A. Yung et al., 2024). Simulations are well converged up to $z\sim 8$ but their predictions spread at $z>8$. However, observational constraints are also scattered at $z>10$, and it is thus still difficult to decisively conclude one physical model is favored over others. The large scatter at $z>10$ in observational data points seems not only due to the Poisson error on small sample size but also to the higher CVs and/or low-$z$ interloper contamination, which will be discussed further in the following section.
@@
Figure 4: The redshift evolution of cosmic UV luminosity density. Black solid line is the median estimation from CANUCS+TEC data in this work, based on the redshift-evolving UV LF measurement in Sec. 4.1, and red shaded area shows the 16th- to 84th-percentile range. Black dashed line is the $\rho_{\rm UV}$ measurement from JOF observation (B. Robertson et al., 2024), which uses similarly rich NIRCam MB+BB filters in high-$z$ galaxy selection. For comparison, literature $\rho_{\rm UV}$ measurements from other JWST observations are shown filled small symbols (P. G. Pérez-González et al., 2023; N. J. Adams et al., 2024; C. T. Donnan et al., 2024; C. J. Willott et al., 2024; Y. Harikane et al., 2025). Gray dash-dot curve marks the prediction of $\rho_{\rm UV}$ evolution assuming a constant star-formation efficiency at $z>6$ by Y. Harikane et al. (2022). Dotted curves present predictions by several theoretical simulations (THESAN-ZOOM, R. Kannan et al. 2025; Universe Machine, P. Behroozi et al. 2020; FLARES, A. P. Vijayan et al. 2021; S. M. Wilkins et al. 2023; SC SAM, L. Y. A. Yung et al. 2024). The $\rho_{\rm UV}$ evolution slope in our work at $z\gtrsim 10$ (black solid line) is somewhat less steep than previous measurements at $z\sim 8$ and in a great agreement with that from JOF data, and starts to deviate from the constant SFE prediction above $z\sim 11$.
@@ 5.1 Cosmic variance
A component of observational uncertainty in measuring the high-$z$ galaxy abundance is the CV (e.g., C. L. Steinhardt et al., 2021; G. Desprez et al., 2024; C. J. Willott et al., 2024; C. K. Jespersen et al., 2025). As shown in Sec. 4.1 and 4.3, UV LFs and $\rho_{\rm UV}$ measurements in our CANUCS+TEC fields are $\sim$0.6 dex lower than in the JOF by B. Robertson et al. (2024), which can be interpreted as the effect of Poisson + CV uncertainties.
The effect of CV appears larger than nominal predictions based on semi-analytical simulations . There are nine galaxies found at $1611.5<z<$ 16 in the JOF, and the predicted uncertainty due to CV on the JOF number count is 23 % (1-sigma), slightly smaller than the Poisson uncertainty (M. Trenti & M. Stiavelli, 2008). The combined fractional error would be 41 %, which leads to $\sim 0.2$ dex uncertainty in the number density estimation including UV LFs, though, our observation shows galaxy abundance measurements can actually be affected by as much as 0.6 dex at this high redshift. Galaxy number statistics at high redshift can thus be considerably affected by field-to-field variance, more severely than the nominal semi-analytical models’ predictions, which is presumably due to bursty star-formation (e.g., G. Sun et al., 2023) or non-universal IMFs (e.g., L. Y. A. Yung et al., 2024) in the early universe that can enhance the field-to-field variance. Therefore, building a larger sample of robust high-$z$ galaxies over wide areas in multiple lines of sight is essential to ascertain the earliest galaxy evolution at $z>10$.
@@ 5.2 Low-$z$ interlopers when selecting $z\gtrsim 10$ galaxies only with Broad Bands
The BB-only catalog selection high-$z$ sample consists of 10 galaxies, four of them are common with our main high-$z$ sample from the full MB+BB photometry (Sec. 3), while six of them are not. The six galaxies are selected as high-$z$ when only BB filter photometry is available, but turn out to be lower-$z$ with the full MB+BB data (”False Positives”). The False Positives, except for CANUCS-2225353, appear as good-looking dropouts with negligible low-$z$ solution probability (less than $\sim$5 %) when the BB-only catalog is used, so it seems inevitable to select them as high-$z$ from BB-only selection (see Appendix A for details of the False Positives). CANUCS-2225353 does pass all selection criteria using the BB-only catalog and is selected as a $z\sim 16$ candidate with F200W-dropout, but the source is also detected in F150W (which is not included as a ”blue” filter in the criteria). In the following computation of the BB-only UV LF, we therefore exclude CANUCS-2225353 and use the remaining nine galaxies as the high-$z$ sample from the BB-only catalog. The comparison of the BB-only selection high-$z$ sample with our main high-$z$ sample indicates that as much as half of $z\gtrsim 10$ galaxy candidates could actually be low-$z$ interlopers when only BB photometry is available. Of course this fraction largely depends on the selection criteria, data quality (e.g. depth), or observation line of sights (not only clustering of $z>10$ galaxies but also lower-$z$ interloper clustering can affect this fraction), but this fraction is roughly consistent with simulations using mock galaxy catalogs by N. J. Adams et al. (2025). Thus the low-$z$ interloper contamination is not negligible, particularly at $z\gtrsim 10$ when only BB photometry data is available.
@@
We then investigate the effect of the low-$z$ interloper contaminants on the UV LF measurements. We follow the same procedure as Sec. 4 on the BB-only catalog selection high-$z$ sample and derive the best estimate UV LF redshift evolution from it. In this BB-only UV LF measurement, we also perform the simulation as we do in Sec. 3.2 with removing the MBs to obtain the completeness of the BB-only catalog selection of $169.5<z<$ 16 galaxies. The completeness gets lower in general with the BB-only catalog selection as compared to the standard high-$z$ selection with MB+BBs.
Figure 5 compares the estimated $\phi_{\star}(z)$ evolution from the main high-$z$ galaxy sample (Sec. 4.1) and that from the BB-only UV LF measurement. The overestimation of the UV LF normalization parameter is severe at higher redshift and it reaches up to $\sim 0.6$ dex at $z\sim 15$. The overestimation of $\phi_{\star}(z)$ similarly affects the $\rho_{\rm UV}(z)$ measurement.The effect of low-$z$ interlopers is particularly critical at $z\gtrsim 15$, because the galaxy abundance at $z\gtrsim 15$ becomes very low while the corresponding lower-$z$ range of the interlopers enters intermediate-$z$ range of $z\sim 4$, where extreme line emitters start to commonly populate (e.g., S. Withers et al., 2023; D. J. Eisenstein et al., 2023). Indeed, the two brightest galaxy candidates in the BB-only catalog selection are both False Positive misidentified as $z\sim 16$ galaxies brighter than $M_{\rm UV}=-20$ mag, similar to CEERS-9331 (P. Arrabal Haro et al., 2023). Low-$z$ interlopers can thus bias the UV LF and UV luminosity density measurements by $\sim 0.6$ dex in BB-only studies. The brighter magnitude bins at higher-$z$ are more likely to be contaminated, and thus LF bright-end studies at $z\gtrsim 10$ need particular caution.
@@ 5.3 Effects of interlopers on other galaxy statistics
The considerable contamination by low-$z$ interlopers in the $z\gtrsim 10$ galaxy sample from the BB-only catalog raises caution not only in UV LFs but also in other statistical studies of earliest galaxy properties. In the following, we compare the False Positives and the robust high-$z$ galaxies identified with MB+BB observations in several aspects to reveal the potential effects of low-$z$ interloper contamination in high-$z$ galaxy studies at $z\gtrsim 10$. We here also utilize the JOF $z>10$ galaxy sample (B. Robertson et al., 2024) as an additional robust high-$z$ galaxy sample, since they select the sample based on similarly rich NIRCam MB+BB filters. The median redshift of CANUCS+TEC+JOF sample is $\langle z\rangle=11.5$.
@@ 5.3.1 Size-$M_{\rm UV}$ relation
Figure 6: Top: cutouts of four examples of False Positives in the F277W filter. They are $1.\!\!^{\prime\prime}6$ on the side. Middle: cutouts of four True Positives in the F277W filter. Bottom: distribution of high-$z$ galaxy candidates in the size-$M_{\rm UV}$ plane. High-$z$ candidates selected with extensive MB+BB filters (red squares, this work; black crosses, JOF) are systematically smaller than a previous calibration of the size-$M_{\rm UV}$ relation at $z\sim 12$ by T. Morishita et al. (2024, black solid line). On the other hand, the False Positives (blue squares) appear systematically larger than the main high-$z$ and JOF candidates and would align well with the size-$M_{\rm UV}$ relation by T. Morishita et al. (2024). The bottom sub-panel shows the systematic offset in $\log(R_{\rm eff})$ of the main high-$z$ candidates from the predicted size by T. Morishita et al. (2024) relation. The median offset is $-0.6$ dex (red line).
The size-$M_{\rm UV}$ relation is one of the key scaling relations of galaxies and it is an a priori assumption in deriving the completeness for the galaxy luminosity function. Figure 6 shows the distribution of the robust high-$z$ galaxy sample from MB+BB data in the size-$M_{\rm UV}$ plane (red squares and black crosses; main sample in this work and from B. Robertson et al. 2024, respectively). The robust high-$z$ galaxy sample distribution is systematically smaller than the size-$M_{\rm UV}$ relation at $1311<z<$ 13 previously calibrated by T. Morishita et al. (2024, black line in the figure; the shaded area presents the 1.5-sigma scatter). This systematic offset cannot be fully explained solely by incompleteness: our sample is $50\%$ complete up to $\log(R_{\rm eff}/{\rm kpc})\sim-0.5$ at $M_{\rm UV}=-19.0$ mag (gray dashed line in the figure), which corresponds to the size-$M_{\rm UV}$ relation median of T. Morishita et al. (2024), while almost all the robust high-$z$ galaxy sample lies well below the T. Morishita et al. (2024) relation. The systematic offset is $\Delta\log(R_{\rm eff})=-0.6$ dex (red line in the bottom panel). On the other hand, if we measure the size and absolute UV magnitude of the False Positives based on the BB-only photometry catalog and photo-$z$ estimation from the BB-only data, they distribute very well along the previously calibrated relation (blue squares). Figure 6 thus demonstrates that the False Positives typically appear larger and brighter than the real high-$z$ galaxies, and the size-$M_{\rm UV}$ relation can be biased towards larger sizes when selecting high-$z$ candidates based on BB-only data. In the completeness simulation (Sec. 3.2), we thus assume a modified size-$M_{\rm UV}$ relation of T. Morishita et al. (2024) offset 0.6 dex smaller. The good agreement between the estimated completeness with this modification and the observed high-$z$ galaxies in the $M_{\rm UV}$-$z$ plane validates this offset.

### v2

@@ 4.3 Cosmic UV Luminosity Density
Given the parameterized UV LF are measured as a function of redshift directly from the data in Sec. 4.1, we can directly compute the redshift evolution of UV luminosity density ($\rho_{\rm UV}(z)$) from the UV LF measurement. We integrate the LF in the Schechter functional form (Eqn. 2) down to $M_{\rm UV}=-17$ mag, and obtain the marginalized $\rho_{\rm UV}(z)$ estimation from the MCMC sampling in Sec. 4.1. Figure 5 shows the median estimation (black solid) with the 16th-84th percentile (red shaded) from our results, comparing with the literature (Perez_Gonzalez2023; Perez_Gonzalez2025; Adams2024ApJ; Donnan2024; Robertson2024ApJ; Willott2024ApJ; Harikane2025ApJ). The overall $\rho_{\rm UV}$ value in this work is somewhat lower than most JWST studies, and in line with Adams2024ApJ and Willott2024ApJ. Some literature claims that significant modifications of dust, IMF, or star formation efficiencies (SFEs) in model assumptions are required based on the overabundance of UV bright galaxies at $z>10$ (e.g., Finkelstein2024ApJ; Harikane2025ApJ; Perez_Gonzalez2025). However, our result together with Adams2024ApJ and Willott2024ApJ should suggest that such significant modifications are not necessarily required.
@@
The log-linear rate of the UV LF evolution $\eta$ is directly connected to the $\rho_{\rm UV}(z)$ redshift evolution, and our result indicates $\rho_{\rm UV}(z)$ decline as $d\log(\rho_{\rm UV})/dz=-0.21^{+0.11}_{-0.12}$ at $10<z<16$. This is less steep than measurements at $z\sim 8-11$ by Adams2024ApJ; Willott2024ApJ and consistent with Robertson2024ApJ at z=1116z=11-16. Importantly, this is also less steep than the prediction by the simple constant extrapolation of the SFE from lower redshift (e.g., Harikane2022, the gray dash-dot line in Figure 5). These results suggest $\rho_{\rm UV}$ evolution gets shallower above $z\sim 11$, and one of the possible interpretations is the SFE may start to deviate at $z>11$ from the simple constant extrapolation from lower redshift.
@@
Figure 5 also compares observations to predictions by several simulations: THESAN-ZOOM (Kannan2025), Universe Machine (Behroozi2020), FLARES (Vijayan2021; Wilkins2023), and SC SAM (Yung2024). Simulations are well converged up to $z\sim 8$ but their predictions spread at $z>8$. However, observational constraints are also scattered at $z>10$, and it is thus still difficult to decisively conclude one physical model is favored over others. The large scatter at $z>10$ in observational data points seems not only due to the Poisson error on small sample size but also to the higher CVs and/or low-$z$ interloper contamination, which will be discussed further in the following section.
@@
Figure 5: The redshift evolution of cosmic UV luminosity density. Black solid line is the median estimation from CANUCS+TEC data in this work, based on the redshift-evolving UV LF measurement in Sec. 4.1, and red shaded area shows the 16th- to 84th-percentile range. Black dashed line is the $\rho_{\rm UV}$ measurement from JOF observation (Robertson2024ApJ), which uses similarly rich NIRCam MB+BB filters in high-$z$ galaxy selection. For comparison, literature $\rho_{\rm UV}$ measurements from other JWST observations are shown filled small symbols (Perez_Gonzalez2023; Adams2024ApJ; Donnan2024; Willott2024ApJ; Harikane2025ApJ). Gray dash-dot curve marks the prediction of $\rho_{\rm UV}$ evolution assuming a constant star-formation efficiency at $z>6$ by Harikane2022. Dotted curves present predictions by several theoretical simulations (THESAN-ZOOM, Kannan2025; Universe Machine, Behroozi2020; FLARES, Vijayan2021; Wilkins2023; SC SAM, Yung2024). The $\rho_{\rm UV}$ evolution slope in our work at $z\gtrsim 10$ (black solid line) is somewhat less steep than previous measurements at $z\sim 8$ and in a great agreement with that from JOF data, and starts to deviate from the constant SFE prediction above $z\sim 11$.
@@ 5.1 Cosmic variance
A component of observational uncertainty in measuring the high-$z$ galaxy abundance is the CV (e.g., Steinhardt2021; Desprez2024; Willott2024ApJ; Jespersen2025ApJ). As shown in Sec. 4.1 and 4.3, UV LFs and $\rho_{\rm UV}$ measurements in our CANUCS+TEC fields are $\sim$0.6 dex lower than in the JOF by Robertson2024ApJ, which can be interpreted as the effect of Poisson + CV uncertainties.
The effect of CV appears larger than nominal predictions based on semi-analytical simulations . There are nine galaxies found at $11.5<z<16$ in the JOF, and the predicted uncertainty due to CV on the JOF number count is 23 % (1-sigma), slightly smaller than the Poisson uncertainty (Trenti2008). The combined fractional error would be 41 %, which leads to $\sim 0.2$ dex uncertainty in the number density estimation including UV LFs, though, our observation shows galaxy abundance measurements can actually be affected by as much as 0.6 dex at this high redshift. Galaxy number statistics at high redshift can thus be considerably affected by field-to-field variance, more severely than the nominal semi-analytical models’ predictions, which is presumably due to bursty star-formation (e.g., Sun2023ApJ) or non-universal IMFs (e.g., Yung2024) in the early universe that can enhance the field-to-field variance. Therefore, building a larger sample of robust high-$z$ galaxies over wide areas in multiple lines of sight is essential to ascertain the earliest galaxy evolution at $z>10$.
@@ 5.2 Low-$z$ interlopers when selecting $z\gtrsim 10$ galaxies only with Broad Bands
The BB-only catalog selection high-$z$ sample consists of 10 galaxies, four of them are common with our main high-$z$ sample from the full MB+BB photometry (Sec. 3), while six of them are not. The six galaxies are selected as high-$z$ when only BB filter photometry is available, but turn out to be lower-$z$ with the full MB+BB data (”False Positives”). The False Positives, except for CANUCS-2225353, appear as good-looking dropouts with negligible low-$z$ solution probability (less than $\sim$5 %) when the BB-only catalog is used, so it seems inevitable to select them as high-$z$ from BB-only selection (see Appendix A for details of the False Positives). CANUCS-2225353 does pass all selection criteria using the BB-only catalog and is selected as a $z\sim 16$ candidate with F200W-dropout, but the source is also detected in F150W (which is not included as a ”blue” filter in the criteria). In the following computation of the BB-only UV LF, we therefore exclude CANUCS-2225353 and use the remaining nine galaxies as the high-$z$ sample from the BB-only catalog. The comparison of the BB-only selection high-$z$ sample with our main high-$z$ sample indicates that as much as half of $z\gtrsim 10$ galaxy candidates could actually be low-$z$ interlopers when only BB photometry is available. Of course this fraction largely depends on the selection criteria, data quality (e.g. depth), or observation line of sights (not only clustering of $z>10$ galaxies but also lower-$z$ interloper clustering can affect this fraction), but this fraction is roughly consistent with simulations using mock galaxy catalogs by Adams2025. Thus the low-$z$ interloper contamination is not negligible, particularly at $z\gtrsim 10$ when only BB photometry data is available.
@@
We then investigate the effect of the low-$z$ interloper contaminants on the UV LF measurements. We follow the same procedure as Sec. 4 on the BB-only catalog selection high-$z$ sample and derive the best estimate UV LF redshift evolution from it. In this BB-only UV LF measurement, we also perform the simulation as we do in Sec. 3.2 with removing the MBs to obtain the completeness of the BB-only catalog selection of $9.5<z<16$ galaxies. The completeness gets lower in general with the BB-only catalog selection as compared to the standard high-$z$ selection with MB+BBs.
Figure 6 compares the estimated $\phi_{\star}(z)$ evolution from the main high-$z$ galaxy sample (Sec. 4.1) and that from the BB-only UV LF measurement. The overestimation of the UV LF normalization parameter is severe at higher redshift and it reaches up to $\sim 0.6$ dex at $z\sim 15$. The overestimation of $\phi_{\star}(z)$ similarly affects the $\rho_{\rm UV}(z)$ measurement.The effect of low-$z$ interlopers is particularly critical at $z\gtrsim 15$, because the galaxy abundance at $z\gtrsim 15$ becomes very low while the corresponding lower-$z$ range of the interlopers enters intermediate-$z$ range of $z\sim 4$, where extreme line emitters start to commonly populate (e.g., Withers2023; Eisenstein2023arXiv). Indeed, the two brightest galaxy candidates in the BB-only catalog selection are both False Positive misidentified as $z\sim 16$ galaxies brighter than $M_{\rm UV}=-20$ mag, similar to CEERS-9331 (Arrabal_Haro2023). Low-$z$ interlopers can thus bias the UV LF and UV luminosity density measurements by $\sim 0.6$ dex in BB-only studies. The brighter magnitude bins at higher-$z$ are more likely to be contaminated, and thus LF bright-end studies at $z\gtrsim 10$ need particular caution.
@@ 5.3 Effects of interlopers on other galaxy statistics
The considerable contamination by low-$z$ interlopers in the $z\gtrsim 10$ galaxy sample from the BB-only catalog raises caution not only in UV LFs but also in other statistical studies of earliest galaxy properties. In the following, we compare the False Positives and the robust high-$z$ galaxies identified with MB+BB observations in several aspects to reveal the potential effects of low-$z$ interloper contamination in high-$z$ galaxy studies at $z\gtrsim 10$. We here also utilize the JOF $z>10$ galaxy sample (Robertson2024ApJ) as an additional robust high-$z$ galaxy sample, since they select the sample based on similarly rich NIRCam MB+BB filters. The median redshift of CANUCS+TEC+JOF sample is $\langle z\rangle=11.5$.
@@ 5.3.1 Size-$M_{\rm UV}$ relation
Figure 7: Top: cutouts of four examples of False Positives in the F277W filter. They are $1.\!\!^{\prime\prime}6$ on the side. Middle: cutouts of four True Positives in the F277W filter. Bottom: distribution of high-$z$ galaxy candidates in the size-$M_{\rm UV}$ plane. High-$z$ candidates selected with extensive MB+BB filters (red squares, this work; black crosses, JOF) are systematically smaller than a previous calibration of the size-$M_{\rm UV}$ relation at $z\sim 12$ by Morishita2024. On the other hand, the False Positives (blue squares) appear systematically larger than the main high-$z$ and JOF candidates and would align well with the size-$M_{\rm UV}$ relation by Morishita2024. The bottom sub-panel shows the systematic offset in $\log(R_{\rm eff})$ of the main high-$z$ candidates from the predicted size by Morishita2024 relation. The median offset is 0.6-0.6 dex (red line).
The size-$M_{\rm UV}$ relation is one of the key scaling relations of galaxies and it is an a priori assumption in deriving the completeness for the galaxy luminosity function. Figure 7 shows the distribution of the robust high-$z$ galaxy sample from MB+BB data in the size-$M_{\rm UV}$ plane (red squares and black crosses; main sample in this work and from Robertson2024ApJ, respectively). The robust high-$z$ galaxy sample distribution is systematically smaller than the size-$M_{\rm UV}$ relation at $11<z<13$ previously calibrated by Morishita2024. This systematic offset cannot be fully explained solely by incompleteness: our sample is $50\%$ complete up to $\log(R_{\rm eff}/{\rm kpc})\sim-0.5$ at $M_{\rm UV}=-19.0$ mag (gray dashed line in the figure), which corresponds to the size-$M_{\rm UV}$ relation median of Morishita2024, while almost all the robust high-$z$ galaxy sample lies well below the Morishita2024 relation. The systematic offset is $\Delta\log(R_{\rm eff})=-0.6$ dex (red line in the bottom panel). On the other hand, if we measure the size and absolute UV magnitude of the False Positives based on the BB-only photometry catalog and photo-$z$ estimation from the BB-only data, they distribute very well along the previously calibrated relation (blue squares). Figure 7 thus demonstrates that the False Positives typically appear larger and brighter than the real high-$z$ galaxies, and the size-$M_{\rm UV}$ relation can be biased towards larger sizes when selecting high-$z$ candidates based on BB-only data. In the completeness simulation (Sec. 3.2), we thus assume a modified size-$M_{\rm UV}$ relation of Morishita2024 offset 0.6 dex smaller. The good agreement between the estimated completeness with this modification and the observed high-$z$ galaxies in the $M_{\rm UV}$-$z$ plane validates this offset.